3. Given the following information about a showroom reno- vation, which activities are on time, which are early, and which are behind schedule? Budgeted Actual Critical Activity A: Plan changes B: Solicit bids C: Select contractor D: Schedule date E: Start renovation Cost $60 $25 $45 $20 $50 Cost $40 S50 S30 $20 S50 Ratio 1.0 0.5 1.5 1.5 0.67
Answer #1

Critical ratio is calculated as (actual progress/schedules progress)*(budgeted cost/actual cost)
Critical ratio is the product of two separate ratios.

Activity A: Plan changes
Budgeted cost=$60
Actual cost=$40
Critical ratio =1
So, (actual progress/scheduled progress)*(budgeted cost/actual cost)=1
=>(actual progress/scheduled progress)*60/40=1
=>(actual progress/scheduled progress)=40/60=.66
This means that actual progress is less than scheduled progress as the value of their ratio is less than 1.

Activity B: Solicit bids
Budgeted cost=$25
Actual cost=$50
Critical ratio =.5
So, (actual progress/scheduled progress)*(budgeted cost/actual cost)=.5
=>(actual progress/scheduled progress)*25/50=.5
=>(actual progress/scheduled progress)=.5*50/25=1
This means that actual progress is same as scheduled progress as the value of their ratio is 1.

Activity C: Select contractor
Budgeted cost=$45
Actual cost=$30
Critical ratio =1.5
So, (actual progress/scheduled progress)*(budgeted cost/actual cost)=1.5
=>(actual progress/scheduled progress)*45/30=1.5
=>(actual progress/scheduled progress)=1.5*30/45=1
This means that actual progress is same as scheduled progress as the value of their ratio is equal to 1.

Activity D: Schedule date
Budgeted cost=$20
Actual cost=$20
Critical ratio =1.5
So, (actual progress/scheduled progress)*(budgeted cost/actual cost)=1.5
=>(actual progress/scheduled progress)*20/20=1.5
=>(actual progress/scheduled progress)=1.5
This means that actual progress is more than scheduled progress as the value of their ratio is more than 1.

Activity E: Start renovation
Budgeted cost=$50
Actual cost=$50
Critical ratio =.67
So, (actual progress/scheduled progress)*(budgeted cost/actual cost)=.67
=>(actual progress/scheduled progress)*50/50=.67
=>(actual progress/scheduled progress)=.67
This means that actual progress is less than scheduled progress as the value of their ratio is less than 1.


Now we need to find the activities that are on time, early and behind schedule.
Activity A: Plan changes
Actual progress is less than scheduled progress, so this activity is behind schedule.
Activity B: Solicit bids
Actual progress is same as scheduled progress, so this activity is on time.
Activity C: Select contractor
Actual progress is same as scheduled progress, so this activity is on time.
Activity D: Schedule date
Actual progress is more than scheduled progress, so this activity is progressing early.
Activity E: Start renovation
Actual progress is less than scheduled progress, so this activity is behind schedule.